OpenStudy (anonymous):

12 m/ min.= _____ cm/ s

OpenStudy (tkhunny):

Multiply by 1, over and over and over... \(\dfrac{12\;meter}{min}\cdot\dfrac{1\;min}{60\;sec}\cdot \dfrac{100\;cm}{1\;meter}\) Watch the units very carefully and simply the expression.